# Elastic Shape Analysis of Movement Data

### v2

@@ 4.3 Nested model comparison for predicting clinical traits
Figure 7 presents a scatterplot of p-values from F-tests of nested models, displayed on a base-10 logarithmic scale. The horizontal axis shows p-values for testing reduced models with full-curve predictors, while the vertical axis represents p-values for reduced models with conventional discrete summary predictors. For readability, the axes are labeled on the original scale.
A red dashed line marks the $\log_{10}(0.05)$ threshold on both axes, corresponding to the $0.05$ significance level. Gray circles represent models where neither reduced model was rejected, a category that includes most clinical traits. This is unsurprising, as many clinical traits have limited association with biomechanics or are inherently difficult to predict, meaning that more complex gait predictors would not be expected to provide much additional value beyond simpler models.
Green plus signs highlight models where the reduced model with full-curve predictors was not rejected, while the reduced model with conventional discrete summary predictors was rejected, meaning conventional summaries provided no additional predictive value. The corresponding traits include Worst knee KLG, Balance confidence, Weight, Waist circumference, Hip circumference, Walking speed, Walk distance, BMI, Male, and Black. These results are consistent with our previous $R^{2}$ comparisons, as these traits were either those for which only full-curve models had significant $R^{2}$ values or those where the $R^{2}$ values were substantially larger than those of models using conventional discrete summaries.
@@
Importantly, there are no cases where the model with full-curve predictors is rejected while the model with conventional discrete summary predictors is not. In fact, no points appear to the left of the vertical red dashed line, indicating that the reduced model with full-curve predictors was never rejected and is consistently a suitable approach.
@@ 5 Conclusion
This paper quantitatively compares the information contained in the data objects of the complete GRF curve versus conventional discrete summaries on disease severity and clinical profiles of patients with OA, demonstrating the greater value of analyzing the full curves. We apply a straightforward nested model comparison to highlight this difference. Furthermore, our shape-based approach offers an intuitive representation of full movement curves, applicable in broader analyses, and reveals insightful modes of variation. Our work is among the first to demonstrate the superior predictive potential of full movement curves over conventional discrete summaries and to compare these representations for predicting disease measures.

### v3

@@ 4.2 Nested Model Comparison
Figure 6 presents a scatterplot of p-values from the LRTs of nested models, displayed on a logarithmic scale. The horizontal axis shows p-values for testing full-curve reduced models, while the vertical axis corresponds to p-values for conventional discrete summary reduced models. For readability, the axes are labeled on the original scale.
The green plus signs highlight traits for which the full-curve reduced model was not rejected, while the conventional discrete summary reduced model was rejected, meaning conventional discrete summaries provided no additional explanatory value. The corresponding traits include joint compressive force; both measures of radiographic OA severity (JSW and KLG); all of the anthropometric and mobility measures; and sex and race. Balance confidence, one of the fall-related traits, was also in this group, though close to the 0.05 threshold.
The gray circles represent cases in which neither reduced model was rejected, which account for roughly half of the traits. These include both inflammatory biomarkers; self-reported pain and function; health-related quality of life measures; all but one fall-related trait; and age. Except for the biomarkers and age, these traits are Likert-type scale scores, which are highly subjective and generally difficult to model. Although inflammatory biomarkers and age are important in OA, there is no known direct relationship with gait biomechanics. Therefore, it is unsurprising that more complex gait variables do not provide significant explanatory value beyond simple discrete summaries for the gray-circle traits.
@@
No points fall to the left of the vertical red dashed line, meaning there were no instances in which the full-curve reduced model was rejected while the conventional discrete summary reduced model was not. Although sex (Male) is near the 0.05 threshold for the full-curve reduced model, it is well below the threshold for the conventional discrete summary reduced model. Overall, these results show that a model based solely on full-curve modes is a consistently suitable approach.
@@ 5 Conclusion
This paper quantitatively demonstrates the extent to which complete GRF curves, compared with conventional discrete summaries, capture information relevant to disease severity and clinical profiles of OA, demonstrating the added value of full-curve analysis. We apply a straightforward nested model comparison to highlight this difference. Furthermore, our shape-based approach illustrates an intuitive representation of full movement curves that is applicable in broader analyses and reveals insightful modes of variation. To our knowledge, this work is among the first to show that analysis of full movement curves yields stronger associations with OA outcomes and OA-related clinical traits than conventional discrete summaries.
@@
Conflict of Interest Statement
The authors report there are no competing interests to declare.
Statement on Generative AI
ChatGPT 4.0 was used for minor coding assistance.